### v1

# Curbing PBHs with PTAs

Abstract
Sizeable primordial curvature perturbations needed to seed a population of primordial black holes (PBHs) will be accompanied by a scalar-induced gravitational wave signal that can be detectable by pulsar timing arrays (PTA). We derive conservative bounds on the amplitude of the scalar power spectrum at the PTA frequencies and estimate the implied constraints on the PBH abundance. We show that only a small fraction of dark matter can consist of stellar mass PBHs when the abundance is calculated using threshold statistics. The strength and the shape of the constraint depend on the shape of the power spectrum and the nature of the non-Gaussianities. We find that constraints on the PBH abundance arise in the mass range $0.1-10^{3}\,M_{\odot}$, with the sub-solar mass range being constrained only for narrow curvature power spectra. These constraints are softened when positive non-Gaussianity is introduced and can be eliminated when $f_{\rm NL}\gtrsim 5$. On the other hand, if the PBH abundance is computed via the theory of peaks, the PTA constraints on PBHs are significantly relaxed, signalling once more the theoretical uncertainties in assessing the PBH abundance. We further discuss how strong positive non-Gaussianites can allow for heavy PBHs to potentially seed supermassive BHs.
I Introduction
Primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974) are currently in the spotlight as they may solve several open questions in astrophysics and cosmology. Different mechanisms and models have been proposed as viable for PBH production and, depending on the formation scenario considered and on the models for PBH formations, a variety of mass functions have been predicted (see Ref. Bagui et al. (2023) for a recent review). Throughout this work, we will assume the standard formation scenario which suggests that PBHs form out of the gravitational collapse after the horizon re-entry of large over-densities in the primordial density contrast field Ivanov et al. (1994); Ivanov (1998); Blinnikov et al. (2016). In such a scenario, a non-negligible PBH abundance requires an enhanced primordial curvature power spectrum at scales smaller than the scale of the Cosmic Microwave Background (CMB), at which the spectral amplitude is around $10^{-9}$ Akrami et al. (2020a).
In addition to producing PBHs, the same primordial curvature perturbations generate tensor perturbations, also known as scalar-induced Gravitational Waves (SIGW). As first shown by Refs. Matarrese et al. (1994); Acquaviva et al. (2003); Mollerach et al. (2004); Ananda et al. (2007); Baumann et al. (2007), when going to second order, scalar and tensor perturbations are coupled and the former provide a source for Gravitational Waves (GW) (see also Refs. Espinosa et al. (2018); Domènech (2021)). The statistical properties of the curvature fluctuations can leave an imprint on this signal. Indeed, the presence of primordial non-Gaussianity (NG) Gangui et al. (1994); Bartolo et al. (2002); Maldacena (2003); Bartolo et al. (2004a, b); Celoria and Matarrese (2020) can modify the spectrum inducing specific features depending on the amplitude of the sourcing scalar power spectrum Garcia-Saenz et al. (2023); Unal (2019); Adshead et al. (2021); Abe et al. (2023); Perna et al. (2024). Moreover, the specific features of the spectra and the NG modifications, as well as the amplitude of the induced SIGW background, depend not only on the shape of the curvature power spectrum considered, but also on the specific scales involved (see, e.g., Cai et al. (2019a, b)). Hence detectable SIGWs offer a potential way to probe early universe cosmology.
The recent Pulsar Timing Array (PTA) data release by the NANOGrav Agazie et al. (2023a, b), EPTA (in combination with InPTA) Antoniadis et al. (2023a, b, c), PPTA Reardon et al. (2023a); Zic et al. (2023); Reardon et al. (2023b) and CPTA Xu et al. (2023) collaborations shows evidence of a Hellings-Downs pattern in the angular correlations which is characteristic of GWs. The most stringent constraints and largest statistical evidence come from the NANOGrav 15-year data (NANOGrav15). Even though it is currently not possible to pinpoint the sources of this signal, this discovery is of extreme importance since it opens a further window into the early universe and has an impact on the GW search at the nHz scale (see, e.g., Ellis et al. (2024); Figueroa et al. (2024)).
Since PTAs are sensitive to frequencies of the order of the nHz, they are associated with the production of PBHs in the stellar mass range Chen et al. (2020); De Luca et al. (2021); Vaskonen and Veermäe (2021); Kohri and Terada (2021); Domènech and Pi (2022); Dandoy et al. (2023); Franciolini et al. (2023a). Crucially, requiring that the produced SIGW background does not exceed the one registered by the NANOGrav15, strongly limits the maximum amplitude of the curvature power spectrum, and thereby the abundance of PBHs in the corresponding mass range. Additionally, if the power spectrum is enhanced at scales larger than accessible by PTAs, PBHs can serve as seeds of supermassive black holes (SMBHs) Duechting (2004); Bernal et al. (2018); Serpico et al. (2020). Nevertheless, in this case, the scales related to the formation process are strongly constrained by CMB spectral distortion analysis of the FIRAS collaboration Fixsen et al. (1996); Chluba et al. (2012a, b); Chluba and Grin (2013); Bianchini and Fabbian (2022) and, without NGs, the allowed PBH abundance is too small to provide a primordial origin for SMBHs seeds.
In this work constraints on PBHs arising from NANOGrav15 are considered in detail with an emphasis on local primordial NGs and assuming different shapes for the primordial power spectrum. After having reviewed the origin of NGs in the standard formation scenario for PBHs in Sec. II, in Sec. III we discuss different formalisms for the computation of the PBH abundance and the role of NGs. In Sec. IV we summarise the computation of the SIGW background. Then, in Sec. V, we derive PTA constraints on the power spectrum and the abundance of PBHs, and comment on how large positive NGs affect the generation of PBHs as seeds for SMBHs. We conclude in Sec. VI.
II Non-Gaussian primordial curvature perturbations
Generally, different models for PBH production require an enhancement of the curvature power spectrum that depends on the model considered. Typical shapes widely used in literature are the log-normal (LN)
$\mathcal{P}_{\zeta}(k)=\frac{{A}}{\sqrt{2\pi}\Delta}\exp\left(-\frac{1}{2\Delta^{2}}\ln^{2}\left(\frac{k}{k_{*}}\right)\right)\,,$ (1)
where $k_{*}$ denotes the peak frequency, or the broken power-law (BPL)
$\mathcal{P}_{\zeta}(k)={A}\,\frac{\left(\alpha+\beta\right)^{\lambda}}{\left[\beta\left(k/k_{*}\right)^{-\alpha/\lambda}+\alpha\left(k/k_{*}\right)^{\beta/\lambda}\right]^{\lambda}},$ (2)
where the parameters $\alpha,\beta>0$ describe the growth and decay of the spectrum around the peak and $\lambda$ the width of the peak. Typically, $\alpha\approx 4$ Byrnes et al. (2019); Karam et al. (2023a) and $\beta\lesssim 4$. The cases considered in this work are reported in Tab. 1.
Case PS shape $f_{\rm NL}$
LN LN                                    $(\Delta=0.5)$ $[0,10]$
BPL1 BPL                  $(\alpha=4,\beta=3,\lambda=1)$ $[0,10]$
BPL2 BPL                  $(\alpha=4,\beta=0.5,\lambda=1)$ $[0,10]$
Table 1: List of benchmark cases considered in this work.
The standard formation scenario assumes that PBHs arise from the gravitational collapse of significant over-densities in the primordial density contrast field Ivanov et al. (1994); Ivanov (1998). Given the random nature of this field, the calculation of PBH abundance is inherently statistical and requires precise knowledge of the probability density function (PDF) of the density fluctuations.
Limiting the analysis to the assumption that the PDF of density fluctuations $\delta$ follows the Gaussian statistics is theoretically flawed for two reasons. First, density fluctuations in the primordial radiation field originate from curvature perturbations $\zeta$, previously stretched on super-horizon scales during inflation, after their horizon re-entry. In the long-wavelength approximation, the equation which relates curvature perturbations to density fluctuations is intrinsically non-linear Harada et al. (2015)
$\delta(\vec{x},t)\!=\!-\frac{2}{3}\frac{\Phi}{(aH)^{2}}e^{-2\zeta(\vec{x})}\bigg{[}\nabla^{2}\zeta(\vec{x})\!+\!\frac{1}{2}\partial_{i}\zeta(\vec{x})\partial_{i}\zeta(\vec{x})\bigg{]},\!\!$ (3)
where $a$ denotes the scale factor, $H$ the Hubble rate, $\Phi$ is related to the equation of state parameter ${w_{\rm u}}$ of the universe. For ${w_{\rm u}}$ constant, $\Phi=3(1+{w_{\rm u}})/(5+3{w_{\rm u}})$ Polnarev and Musco (2007). We dropped the explicit $\vec{x}$ and $t$ dependence for brevity. For this reason, even under the assumption that curvature perturbations follow exact Gaussian statistics, density fluctuations inherit an unavoidable amount of NG from non-linear (NL) corrections De Luca et al. (2019); Young et al. (2019); Germani and Sheth (2020). Second, there is no guarantee that $\zeta$ is a Gaussian field – we refer to such cases as primordial NGs. Local primordial NGs can be described by a relation
$\zeta(\vec{x})=F(\zeta_{\rm G}(\vec{x}))$ (4)
between $\zeta$ and its Gaussian counterpart $\zeta_{G}$, which depends on the mechanism that generates the enhanced power spectrum at small scales (see the related discussion in Ref. Ferrante et al. (2023a)). Such NGs are thus generically independent of large-scale NGs constrained by CMB data (e.g. Akrami et al. (2020b)). Copious PBH production via critical collapse of large overdensities is achievable in a wide range of scenarios including single-field inflation with an inflation point in the inflaton’s potential (usually called USR models) Garcia-Bellido and Ruiz Morales (2017); Pi et al. (2018); Kannike et al. (2017); Ballesteros et al. (2020); Inomata et al. (2017); Iacconi et al. (2022); Kawai and Kim (2021); Bhaumik and Jain (2020); Cheong et al. (2021); Inomata et al. (2018); Dalianis et al. (2019); Motohashi et al. (2020); Hertzberg and Yamada (2018); Ballesteros and Taoso (2018); Karam et al. (2023a); Rasanen and Tomberg (2019); Balaji et al. (2022); Frolovsky and Ketov (2023); Dimopoulos (2017); Germani and Prokopec (2017); Choudhury and Mazumdar (2014); Ragavendra and Sriramkumar (2023); Cheng et al. (2022); Franciolini et al. (2023b); Karam et al. (2023b); Mishra et al. (2023); Cole et al. (2023); Karam et al. (2023b); Frosina and Urbano (2023); Franciolini and Urbano (2022); Choudhury et al. (2024); Wang et al. (2024); Stamou (2021, 2024); Heydari and Karami (2022a, b, 2024) and models with spectator field, i.e., the curvaton Enqvist and Sloth (2002); Lyth and Wands (2002); Sloth (2003); Lyth et al. (2003); Dimopoulos et al. (2003); Kohri et al. (2013); Kawasaki et al. (2013a, b); Carr et al. (2017); Ando et al. (2018a, b); Chen and Cai (2019); Liu and Prokopec (2021); Pi and Sasaki (2023); Cai et al. (2021); Liu (2023); Chen et al. (2023); Torrado et al. (2018); Chen et al. (2023); Wilkins and Cable (2024); Inomata et al. (2024).
To be as model-independent as possible we follow the common approach and describe $\zeta$ by the usual power series expansion
$\zeta=\zeta_{\rm G}+\frac{3}{5}f_{\rm NL}(\zeta_{\rm G}^{2}-\langle\zeta_{\rm G}^{2}\rangle)+\frac{9}{25}g_{\rm NL}\zeta_{\rm G}^{3}+\dots\,,$ (5)
where $\zeta_{\rm G}$ obeys the Gaussian statistics while the coefficients $f_{\rm NL}$, $g_{\rm NL}$ encode deviations from the Gaussian limit. In general, the coefficients can depend on the scale of the perturbation. We omit this possibility here.
III PBH formation
A PBH is formed when a sufficiently large primordial overdensity collapses after horizon re-entry. The mass of the resulting PBH follows a critical scaling law
$M_{\rm PBH}=\mathcal{K}M_{k}(\mathcal{C}-\mathcal{C}_{\rm th})^{\gamma}\,,$ (6)
where
$M_{k}\approx 17M_{\odot}\left(\frac{k}{10^{6}{\rm Mpc}^{-1}}\right)^{-2}\,\left(\frac{g_{*}}{10.75}\right)^{-1/6}\,$ (7)
is the mass contained within a Hubble horizon corresponding to a comoving scale $k$. The threshold $\mathcal{C}_{\rm th}$ is estimated following Ref. Musco et al. (2021) and depends on the shape of the power spectrum. We also account for the time dependence of $\mathcal{C}_{\rm th}$, the critical exponent $\gamma$ and the prefactor $\mathcal{K}$ during the QCD phase transition Musco et al. (2024).
The fraction of energy density $\beta_{k}(M_{\rm PBH}){\rm d}\ln M_{\rm PBH}$ collapsing into PBHs can be estimated as
$\!\!\!\beta_{k}(M_{\rm PBH})\!=\!\int_{\mathcal{C}_{\rm th}}^{\infty}\!{\rm d}\mathcal{C}\frac{M_{\rm PBH}(\mathcal{C})}{M_{k}}P_{k}(\mathcal{C})\delta\left[\ln\frac{M_{\rm PBH}}{M_{\rm PBH}(\mathcal{C})}\right]\!,$ (8)
where $P_{k}(\mathcal{C})$ denotes the probability a BH will form in the Hubble patch. The PBH mass function can be obtained directly from the collapse fraction:
$\!\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}$ $\!=\!\frac{1}{\Omega_{\rm DM}}\int\frac{{\rm d}M_{k}}{M_{k}}\,\beta_{k}(M_{\rm PBH})\left(\frac{M_{\rm eq}}{M_{k}}\right)^{1/2}\!\!,$ (9)
where $M_{\rm eq}\approx 2.8\times 10^{17}\,\,M_{\odot}$ is the horizon mass at the time of matter-radiation equality and $\Omega_{\rm DM}=0.12h^{-2}$ is the cold dark matter density Akrami et al. (2020a). To characterise the PBH population, we will consider the PBH abundance and the mean PBH mass,
$f_{\rm PBH}$ $=\int\frac{{\rm d}M_{\rm PBH}}{M_{\rm PBH}}\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}\,,$ (10)
$\langle M_{\rm PBH}\rangle$ $=f_{\rm PBH}\left(\int\frac{{\rm d}M_{\rm PBH}}{M_{\rm PBH}^{2}}\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}\right)^{-1}\,,$
where the mean PBH mass is computed with respect to the PBH number density (see e.g. Andrés-Carcasona et al. (2024)). This offers a clear basis for comparing constraints on PBHs with extended mass functions.
The above description can be applied to estimate PBH abundance using both the threshold statistics and peaks theory. However, the shape and the origin of $P_{k}(\mathcal{C})$ in these two approaches are different.
Threshold statistics
dictates that the PBH formation probability can be estimated from the statistics of the compaction function $\mathcal{C}$ Ferrante et al. (2023a); Gow et al. (2023), generically defined as twice the local mass excess over the areal radius [^1] Using the threshold statistics approach, which relies on average compaction profiles Ianniccari et al. (2024), corrections to the horizon crossing and from the non-linear radiation transfer function Franciolini et al. (2023c); De Luca et al. (2023) can affect the PBH abundance. Since the amount of these corrections is still not well understood we leave their inclusions for a future work.. The latter can be estimated in scenarios with non-Gaussian curvature perturbations $\zeta$ in case they can be characterized via an auxiliary Gaussian field $\zeta_{G}$ as outlined in Sec. II. The compaction function $\mathcal{C}=\mathcal{C}_{1}-\mathcal{C}_{1}^{2}/(4\Phi)$ can then be constructed from $\mathcal{C}_{1}=\mathcal{C}_{\rm G}\,{\rm d}F/{\rm d}\zeta_{\rm G}$ which can be expressed in terms of the Gaussian field $\mathcal{C}_{\rm G}=-2\Phi\,r\,\zeta_{\rm G}^{\prime}$.
The PBH formation probability in a Hubble patch is then
$P_{k}(\mathcal{C})=\int_{\mathcal{D}}\delta(\mathcal{C}-\mathcal{C}(\mathcal{C}_{\rm G},\zeta_{\rm G}))P_{{\rm G},k}(\mathcal{C}_{\rm G},\zeta_{\rm G}){\rm d}\mathcal{C}_{\rm G}{\rm d}\zeta_{\rm G}\,,$ (11)
where the domain of integration $\mathcal{D}=\left\{\mathcal{C}(\mathcal{C}_{\rm G},\zeta_{\rm G})>\mathcal{C}_{\rm th}~{}\land~{}\mathcal{C}_{1}(\mathcal{C}_{\rm G},\zeta_{\rm G})<2\Phi\right\}$. and the Gaussian components are distributed as
$P_{\mathrm{G}}\left(\mathcal{C}_{\mathrm{G}},\zeta_{\mathrm{G}}\right)=\frac{e^{\left[-\frac{1}{2\left(1-\gamma_{cr}^{2}\right)}\left(\frac{\mathcal{C}_{\mathrm{G}}}{\sigma_{c}}-\frac{\gamma_{cr}\zeta_{\mathrm{G}}}{\sigma_{r}}\right)^{2}\!-\!\frac{\zeta_{\mathrm{G}}^{2}}{2\sigma_{r}^{2}}\right]}}{2\pi\sigma_{c}\sigma_{\tau}\sqrt{1-\gamma_{cr}^{2}}}.$ (12)
The correlators are given by
$\sigma_{c}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{4}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (13a)
$\sigma_{cr}^{2}=\frac{2\Phi}{3}\!\int_{0}^{\infty}\!\!\frac{{\rm d}k}{k}\!\left(kr_{m}\right)^{2}\!W\!\!\left(k,r_{m}\right)\!W_{s}\!\left(k,r_{m}\right)\!P^{T}_{\zeta}\!\,,$ (13b)
$\sigma_{r}^{2}=\int_{0}^{\infty}\frac{{\rm d}k}{k}W_{s}^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,,$ (13c)
with $P^{T}_{\zeta}=T^{2}\left(k,r_{m}\right)P_{\zeta}(k)$, and $\gamma_{cr}\equiv\sigma_{cr}^{2}/\sigma_{c}\sigma_{\tau}$. We have defined $W\left(k,r_{m}\right),$ $W_{s}\left(k,r_{m}\right)$ and $T\left(k,r_{m}\right)$ as the top-hat window function, the spherical-shell window function, and the radiation transfer function Young (2022).
The PBH mass function (9) can be expressed as
$\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}\!=\!\frac{1}{\Omega_{\rm DM}}\int_{M_{\rm H}^{\rm min}}\frac{dM_{\rm H}}{M_{\rm H}}\left(\frac{M_{\rm eq}}{M_{\rm H}}\right)^{1/2}\left(\frac{M_{\rm PBH}}{\mathcal{K}M_{\rm H}}\right)^{\frac{1+\gamma}{\gamma}}$
×$\times\frac{{\cal K}}{\gamma\sqrt{\Lambda}}\int{\rm d}\zeta_{{\rm G},k}P_{\rm G}({\cal C}_{\rm G}(M_{\rm PBH},\zeta_{\rm G}),\zeta_{\rm G})\left(\frac{dF}{d\zeta_{\rm G}}\right)^{-1}\!\!\!\!,$ (14)
where $\Lambda=1-({\cal C}_{\rm th}-\left(M_{\rm PBH}/(\mathcal{K}M_{H})\right)^{1/\gamma})/\Phi$.
Theory of peaks
It is well known that the Press-Schechter approach does not agree with the theory of peaks (see, e.g., Refs. Green et al. (2004); Young et al. (2014); De Luca et al. (2019)). A technical drawback of peaks theory is that it is not clear how to include NGs in the computation of the abundance in this approach using a generic functional form for the curvature perturbation field as in Eq. (4). Hence, for comparison, we limit our analysis for the peak theory approach to the case of negligible primordial NGs [^2] The impact of local-type NG on PBH abundance in peak theory is expected to be reduced, especially for higher-order terms Young (2022) and for peaked power spectra. Indeed, the compaction is volume-averaged over the scale of the perturbation and depends on the curvature perturbation at its edge, rather than its centre..
In this approach, the starting point is the number density of peaks with a height in the interval $(\nu,\nu+{\rm d}\nu)$. When $\nu\gg 1$, it is given by Bardeen et al. (1986)
$\frac{{\rm d}\mathcal{N}}{{\rm d}\nu}=\frac{1}{4\pi^{2}r_{m}^{3}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)\,,$ (15)
where we introduced the rescaled peak height $\nu\equiv\mathcal{C}_{1}/\sigma_{c}$ and the first rescaled moment of the distribution Balaji et al. (2023)
$\sigma_{cc}^{2}=\frac{4\Phi^{2}}{9}\int_{0}^{\infty}\frac{{\rm d}k}{k}\left(kr_{m}\right)^{6}W^{2}\left(k,r_{m}\right)P^{T}_{\zeta}\,.$ (16)
The number of peaks within a Hubble volume is therefore
$\frac{{\rm d}N_{k}}{{\rm d}\nu}=\frac{4\pi}{3}r_{m}^{3}\mathcal{N}$ (17)
and, since $N_{k}\ll 1$, the probability of finding at least one peak within a range $(\nu,\nu+{\rm d}\nu)$ within a Hubble volume is $P_{k}\approx{\rm d}N_{k}$. The mass fraction of PBHs is then Yoo et al. (2018, 2019); Gow et al. (2021); Franciolini et al. (2022)
$P_{k}(\mathcal{C})=\frac{1}{3\pi}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}\nu^{3}\exp\left(-\frac{\nu^{2}}{2}\right)$ (18)
and the domain of integration agrees with Eq. 8. Thus, peaks theory predicts an additional $\nu^{3}$ factor to the collapse probability in the Gaussian case.
Recasting the quantities in terms of $M_{k}$, as in (III), we can express the PBH abundance as
$\frac{{\rm d}f_{\rm PBH}}{{\rm d}\ln M_{\rm PBH}}=$ $\frac{1}{\Omega_{\rm DM}}\int_{M_{\rm H}^{\rm min}}\frac{dM_{\rm H}}{M_{\rm H}}\left(\frac{M_{\rm eq}}{M_{\rm H}}\right)^{1/2}\left(\frac{M_{\rm PBH}}{\mathcal{K}M_{\rm H}}\right)^{\frac{1+\gamma}{\gamma}}$
×$\times\frac{{\cal K}}{\gamma}\frac{\left(2\Phi\left(1-\sqrt{\Lambda}\right)\right)^{3}}{3\pi\sigma_{c}^{4}\Lambda^{1/2}}\left(\frac{\sigma_{cc}}{\sigma_{c}}\right)^{3}$
×$\times\exp\left[{\frac{-2\Phi^{2}}{\sigma_{c}^{2}}\left(1-\sqrt{\Lambda}\right)^{2}}\right]\,.$ (19)
IV Scalar-induced gravitational waves
Primordial curvature fluctuations large enough to generate PBHs will also induce a non-negligible SIGW background. At the second order in perturbation theory, scalar modes can source tensor modes (for a review, see Ref. Domènech (2021)). Starting from Einstein’s equations and omitting anisotropic stress, the equations of motion for GWs in Fourier space are
$h^{\prime\prime}_{\lambda}(\textbf{k},\eta)+2\mathcal{H}h^{\prime}_{\lambda}(\textbf{k},\eta)+k^{2}h_{\lambda}(\textbf{k},\eta)=4\mathcal{S}_{\lambda}(\textbf{k},\eta)\,,$ (20)
with $\lambda$ indicating the polarization and the being the derivative with respect to the conformal time $\eta$. The source term on the right-hand side reads Baumann et al. (2007)
$\mathcal{S}_{\lambda}(\textbf{k},\eta)\!=\!\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})f(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (21)
where $Q$ is a projection factor and $f$ is a function that contains the linear evolution of the Newtonian potential after horizon re-entry and the transfer function mapping them to the curvature perturbation $\zeta$. Eq. (20) can be solved using the Green’s functions approach
$h_{\lambda}(\textbf{k})\!=\!4\int\frac{{\rm d}^{3}\textbf{q}}{(2\pi)^{\frac{3}{2}}}Q_{\lambda}(\textbf{k},\textbf{q})I(|\textbf{k}-\textbf{q}|,q,\eta)\zeta_{\textbf{q}}\zeta_{\textbf{k}-\textbf{q}}\,,$ (22)
where $I$ is the transfer function. To speed up our analysis, we assume perfect radiation domination and do not account for the variation of sound speed during the QCD era (see, for example, Hajkarim and Schaffner-Bielich (2020); Abe et al. (2021)) which also leads to specific imprints in the low-frequency tail of any cosmological SGWB Franciolini et al. (2024). On top of that, cosmic expansion may additionally be affected by unknown physics in the dark sector, which can, e.g., lead to a brief period of matter domination of kination Ferreira and Joyce (1998); Pallis (2006); Redmond et al. (2018); Co et al. (2022); Gouttenoire et al. (2021); Chang and Cui (2022). Explicit analytical expressions for the kernel in different epochs can be found in Domènech (2020), after performing the oscillation average well inside the horizon.
The GW dimensionless power spectrum $\mathcal{P}_{h}(k)$ is defined as
$\langle h_{\lambda_{1}}(\mathbf{k}_{1})h_{\lambda_{2}}(\mathbf{k}_{2})\rangle=\delta^{3}(\mathbf{k}_{1}+\mathbf{k}_{2})\delta_{\lambda_{1}\lambda_{2}}\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{h,\lambda_{1}}(k_{1})$ (23)
and hence
$\Omega_{\rm GW}(k,\eta)=\frac{1}{48}\left(\frac{k}{a(\eta)H(\eta)}\right)\sum_{\lambda=+,\times}\overline{\mathcal{P}_{h,\lambda}(k)}$ (24)
at the production. In the Gaussian case, one obtains
$\Omega_{\rm GW}(k,\eta)|_{\rm G}=$ $\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}\int_{0}^{\infty}{\rm{d}}t\int_{-1}^{1}{\rm{d}}s\hskip 2.84544pt\frac{1}{u^{2}v^{2}}$
×$\times\overline{\tilde{J}^{2}(u,v,x)}\hskip 2.84544pt\mathcal{P}_{\zeta_{g}}\mathcal{P}_{\zeta_{g}}(vk)\mathcal{P}_{\zeta_{g}}(uk)\,,$ (25)
where $x=k\eta$, $t$ and $s$ are dimensionless variables introduced to ease the numerical analysis, $\tilde{J}$ is related to the transfer function $I$ defined before and the overbar indicates an oscillation average after the modes are well inside the horizon. Furthermore $u=(t+s+1)/2$ and $v=(t-s+1)/2$. The spectrum observed today can be obtained using entropy conservation
$h^{2}\Omega_{\rm GW}|_{k,0}=h^{2}\Omega_{\rm rad,0}\left(\frac{g_{s,\eta_{0}}}{g_{s,\eta_{k}}}\right)^{\frac{4}{3}}\frac{g_{\rho,\eta_{k}}}{g_{\rho,\eta_{0}}}\Omega_{\rm GW}(k,\eta)\,,$ (26)
where we used the energy ($g_{\rho}$) and entropy density ($g_{s}$) degrees of freedom of Ref. Borsanyi et al. (2016). According to Ref. Aghanim et al. (2020), we consider $h^{2}\Omega_{\rm rad,0}=4.2\cdot 10^{-5}$.
IV.1 Effects of primordial NG on SIGW
As shown in Kohri and Terada (2018); Cai et al. (2019a, b); Hajkarim and Schaffner-Bielich (2020); Atal and Domènech (2021); Yuan and Huang (2021); Domènech et al. (2022); Garcia-Saenz et al. (2023); Liu et al. (2024); Yuan et al. (2023); Li et al. (2023); Unal (2019); Adshead et al. (2021); Abe et al. (2023); Perna et al. (2024) primordial NGs can leave a non-negligible imprint on the SIGW power spectrum. Hence, considering only the quadratic contribution $f_{\rm NL}$ in the local expansion in Eq. (5) [^3] We specify that, in principle, all the terms in the power series expansion in Eq. (5) should be included. Being interested mainly in the qualitative effect of NGs, we truncated the power series at the quadratic contribution, $f_{\rm NL}$, setting all other terms to 0. These additional NG corrections would further shift the results in Fig. 3., at the next-to-leading order one obtains (see, e.g., Unal (2019); Adshead et al. (2021))
$\Omega_{\rm GW}(k,\eta)|_{\rm t}$ $=\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$
$\times$ $\cos 2\varphi_{12}\frac{u_{1}v_{1}}{(u_{2}v_{2})^{2}}\frac{1}{w_{a,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(u_{2}k)\mathcal{P}_{\zeta_{g}}(w_{a,12}k)\,,$ (27)
$\Omega_{\rm GW}(k,\eta)|_{\rm u}$ $=\hskip 2.84544pt\frac{1}{12\pi}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}\int_{0}^{2\pi}{\rm{d}}\varphi_{12}$
$\times$ $\cos 2\varphi_{12}\frac{u_{1}u_{2}}{(v_{1}v_{2})^{2}}\frac{1}{w_{b,12}^{3}}\overline{\tilde{J}(u_{1},v_{1},x)\tilde{J}(u_{2},v_{2},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(v_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}k)\mathcal{P}_{\zeta_{g}}(w_{b,12}k)$ (28)
$\Omega_{\rm GW}(k,\eta)|_{\rm hyb}$ $=\hskip 2.84544pt\frac{1}{12}\left(\frac{k}{a(\eta)H(\eta)}\right)^{2}f_{\rm NL}^{2}$
$\times$ $\int_{0}^{\infty}{\rm{d}}t_{1}\int_{-1}^{1}{\rm{d}}s_{1}\int_{0}^{\infty}{\rm{d}}t_{2}\int_{-1}^{1}{\rm{d}}s_{2}$
$\times$ $\frac{1}{(u_{1}v_{1}u_{2}v_{2})^{2}}\hskip 2.84544pt\overline{\tilde{J}^{2}(u_{1},v_{1},x)}$
$\times$ $\mathcal{P}_{\zeta_{g}}(u_{1}k)\mathcal{P}_{\zeta_{g}}(v_{2}v_{1}k)\mathcal{P}_{\zeta_{g}}(u_{2}v_{1}k)\,.$ (29)
As in the Gaussian case, the $u_{i}$ and $v_{i}$ are related similarly to the integration variables $t_{i}$ and $s_{i}$.
The GW spectrum can then be expressed as
$\Omega_{{\rm GW}}|_{k,0}$ $=A^{2}\Omega_{\rm GW}|_{{\rm G}}(k/k_{*},0)$ (30)
$+A^{3}f_{\rm NL}^{2}\Omega_{\rm GW}|_{{\rm NG}}(k/k_{*},0)\,,$
[…]
Figure 1: Left panel: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum (1) with $\Delta=0.5$ assuming Gaussian fluctuations and changing the window function. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, assuming NGs arising solely from the non-linear corrections. Right panel: Inferred constraints on PBH abundance.
V Constraints on PBHs
V.1 PTA observations
Consider the constraints on the SIGW signal from the current PTA observations. Using from the NANOGrav 15 year dataset Agazie et al. (2023a); Collaboration (2023), we compute the observational upper bound at 95% confidence level on $h^{2}\Omega^{\rm bound}_{\rm GW}$ in each of the first 14 frequency bins. The frequency of bin ”$i$” is given by $f_{i}=i/T$, where $T=16.03\,{\rm yr}=(1.98\,{\rm nHz})^{-1}$ is the time of observation. We then compare the observation with theoretical prediction $h^{2}\Omega_{\rm GW}$ in Eq. (30) in each frequency bin. The constraint on $h^{2}\Omega_{\rm GW}$ in any given bin implies an upper bound on $A$ for a given model of NGs and a fixed shape of the curvature power spectrum. To combine the constraints of individual bins, we consider the strongest constraint on $A$. The upper bound on $A$ is then mapped to an upper bound on $f_{\rm PBH}$. More specifically, for every fixed benchmark model in Table 1, we can map the pair $(A,k_{*})$ uniquely into the pair $(f_{\rm PBH},\langle M_{\rm PBH}\rangle)$.
Due to the finite observational time, the spectral resolution of the detector is limited and can be estimated to be approximately $\Delta f=1/T=1.98\,{\rm nHz}$. This will limit the PTA sensitivity to sharp features in the SIGW spectrum. To account for this, the contribution to $i$-th bin can be estimated as
$\Omega_{{\rm GW},i}=\int\frac{{\rm d}f}{f}\Omega_{\rm GW}^{\rm th}(f)W(f-f_{i})\,$ (31)
where $W$ denotes a window function. To understand the effect of limited spectral resolution, we will consider three different idealized cases
$W_{\rm D}(f)$ $\propto\delta\left(f\right)\,,$ (32)
$W_{\rm S}(f)$ $\propto{\rm sinc}^{2}\left(\pi\frac{f}{\Delta f}\right)\,,$ (33)
$W_{\rm TH}(f)$ $\propto\theta\left(f+\frac{\Delta f}{2}\right)\theta\left(\frac{\Delta f}{2}-f\right)\,,$ (34)
which correspond to an infinitely accurate frequency resolution, a sharp binning of frequencies and a time-domain top-hat filter on $h_{\rm c}(t)$ with duration $T$, respectively. The window functions are normalized so that $\int{\rm d}\ln f\,W(f-f_{i})=1$. These window functions represent different limiting cases and can thus capture systematics related to the spectral resolution not otherwise accessible in our simplified analysis. As it can be seen from Fig. 1, the effect is expectedly larger at large scales and for heavy PBHs as these are mostly related to the signal at small frequencies. For instance, at $k\simeq 10^{6}\,{\rm{Mpc}}^{-1}$, $W_{\rm D}(f)$ gives a mildly stronger constraint on $A$ when compared to the other choices, as it can resolve sharp peaks in the SIGW spectrum. This is reflected in the constraints on the PBH abundance, as shown on the right panel of Fig. 1. This behaviour is reverted at even smaller $k_{*}$, and $W_{\rm D}(f)$ is less constraining than the other choices. For the remainder of the analysis, we will consider the Dirac window function $W_{\rm D}(f)$.
Figure 2: Left panel: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum, Eq. (1), with $\Delta=0.5$ (blue), and on a broken power-law curvature power spectrum, Eq. (2), with $\alpha=4$, $\gamma=1$ for $\beta=3$ (pink) and $\beta=0.5$ (red), assuming Gaussian fluctuations. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, given the non-linearity-only scenario. Right panel: The implied NANOGrav15 constraints on PBH abundance.
Figure 3: Left panels: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum (1) with $\Delta=0.5$ (top panels) and on a broken power-law curvature power spectrum (2) with $\alpha=4$, $\gamma=1$ for $\beta=3$ (middle panels) and $\beta=0.5$ (bottom panels), assuming quadratic primordial NGs only with $|f_{\rm NL}|\in[0,10]$. The nearly horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics with several $f_{NL}$ values. Right panels: Inferred constraints on the PBH abundance using threshold statistics.
Another important theoretical uncertainty when estimating constraints on the PBH abundance arises due to the choice of the prescription when computing the PBH abundance. The right panels of Figs. 1 and 2 show the constraints on the PBH abundance arising when the abundance is computed using threshold statistics (solid lines) and peaks theory (dashed lines). In both figures, primordial NGs are neglected ($f_{\rm NL}=0$) so the NGs arise only from the non-linear relation between density and curvature perturbations. As one can see, almost the entire parameter space accessible by LVK via GWs from PBH binary mergers Hütsi et al. (2021); Romero-Rodriguez et al. (2022); Franciolini et al. (2022); Andrés-Carcasona et al. (2024) is excluded when threshold statistics is assumed. On the other hand, with the peaks theory, only a small portion of the parameter space relevant to LVK is constrained. Despite this significant theoretical uncertainty, PTA observations can constrain the PBH abundance in both cases.
The dependence of the constraints on the shape of the power spectrum is illustrated in Fig. 2. While relatively narrow spectra, such as the LN and the BPL1 cases, produce similar results, broader spectra, represented by the BPL2 case are constrained also at larger scales due to the broadness of the spectrum. We remark that the constraints at large scales $k_{*}\lesssim 5\times 10^{6}{\rm Mpc}^{-1}$ arise mostly from the first bin by NANOGrav15. This is also the region in which the SIGW are most strongly constrained. In a similar vein, the constraints on the PBH abundance in the threshold statistics case are quite similar for all spectra except at large masses $\langle M_{\rm PBH}\rangle$. In particular, the strongest constraints touch the dashed region in Fig. 2 which corresponds to having less than a single PBH within the current Hubble volume. This line is given by $f_{\rm PBH}4\pi\Omega_{\rm DM}M_{\rm pl}^{2}/H_{0}<\langle M_{\rm PBH}\rangle$.
Fig. 3 shows the effect of primordial NGs. Firstly, we underline that the NG corrections to SIGW depend on $f_{\rm NL}^{2}$ and are thus indifferent to the sign of $f_{\rm NL}$. On the other hand, this latter strongly impacts $f_{\rm PBH}$. Indeed, negative primordial NGs suppress the tail of the PDF with respect to the Gaussian case, that is, they reduce the probability of large fluctuations that cross the threshold. So, a larger amplitude would be needed to obtain the same PBH abundance. Positive primordial NGs, on the other hand, have the opposite effect as they make large fluctuations more likely.
As shown on the left panels of Fig. 3, for small $f_{\rm NL}$ we observe that the corrections to the SIGW affect only amplitudes at the tails, both at smaller and higher scales. A larger positive $f_{\rm NL}$, on the other hand, further shifts the constant $f_{\rm PBH}$ curves downward. As expected, this reduces the allowed amplitude $A$ for a given $k_{*}$. Conversely, for a fixed PBH abundance, positive primordial NGs decrease the required amplitude $A$ with respect to the Gaussian case. Since the latter effect is larger than the reduction of the SIGW strength, positive primordial NGs can completely remove constraints on the PBH abundance.
The situation differs when considering negative NGs. Since negative NGs imply an upward shift in the power spectrum amplitude in order to achieve the same PBH abundance, the constraints become more stringent.
One might mistakenly assume that by increasing the negative value of the coefficient $f_{\rm NL}$, the required power spectral amplitude for the same PBH abundance would always rise, tightening the constraints indefinitely. However, this is not the case. As shown in Ref. Franciolini et al. (2023a), the required amplitude actually reaches a maximum for $f_{\rm NL}\simeq-2$ and then decreases, eventually approaching the amplitude of the Gaussian case for very large negative values of $f_{\rm NL}$. Consequently, the constraints for the case of $f_{\rm NL}=-2$ can be considered as the most stringent.
Nevertheless, this claim should be taken cum grano salis. Indeed, it is important to note that the prescription outlined in Ref. Musco et al. (2021) to compute the threshold for PBH collapse only accounts for NGs arising from the non-linear relation between the density contrast and the curvature perturbations. In principle, also primordial NGs beyond the quadratic approximation should be taken into account when computing the threshold value. Following Refs. Kehagias et al. (2019); Escrivà et al. (2022), it appears that the effect on the threshold is small only for positive NGs and for small negative NGs ($f_{\rm NL}\gtrsim-2$) and at most of the order of a few percent. For more negative NGs, there are huge uncertainties on the exact value of the threshold, which could potentially modify the constraints introduced above [^5] One alternative approach is to use the averaged value over a sphere of radius equal to the location of the maximum of the compaction function, $\bar{C}_{\rm th}=2/5$, since this value is fully non-perturbative and independent of NGs Escrivà et al. (2020); Kehagias et al. (2024). As discussed in Ref. Ianniccari et al. (2024), this requires determining the statistics of the curvature perturbation and computing the connected cumulants, which is a complex task and is left for future work..
For completeness, we analyze a common scenario for PBH formation.
In Ultra slow roll (USR) models of single-field inflation, the peak in $\mathcal{P}_{\zeta}$ arises from a brief phase of ultra-slow-roll followed by constant-roll inflation dual to it Atal and Germani (2019); Biagetti et al. (2018); Karam et al. (2023a). In this case, the NGs can be related to the large $k$ spectral slope Atal et al. (2019); Tomberg (2023),
$\zeta=-\frac{2}{\beta}\log\left(1-\frac{\beta}{2}\zeta_{\rm G}\right).$ (35)
In Fig. 4 we show a similar plot as in Fig. 2 but computing the abundance of PBHs using the full non-gaussian relation reported in Eq.(35) for the two broken power law cases reported in Tab. 1. We compute the abundance using exclusively the method based on the threshold statistics on the compaction function, and as we can see from the plot, with respect to the Gaussian approximation the constraints are highly relaxed and, differently from the Gaussian case, they are of the same order of magnitude. The reason is the following: expanding Eq. (35) at second order, we get $f_{\rm NL}=5\beta/12$, then using the reference points $\beta=3$ and $\beta=0.5$ shown in Fig. 4, one finds $f_{\rm NL}=5/4$ and $f_{\rm NL}=5/24$. As a consequence, the narrower power spectrum, which has stronger constraints in the Gaussian approximation, due to larger primordial non-gaussianities, has constraints similar to the broader case.
Moreover, with respect the impact on the PBH abundance, the NG corrections to the SIGW are irrelevant.
Figure 4: The NANOGrav15 constraints on the PBH abundance using threshold statistics assuming the two cases for the broken power law reported in Tab. 1 assuming the gaussian approximation (solid) and with the full non-gaussian relation in Eq. (35).
[…]
Figure 5: Constraints on the amplitude of the power spectrum, assuming negligible primordial NGs, from the FIRAS experiments, CMB, $\alpha$-Lyman and NANOGrav15 for the same cases reported in Tab. 1.
V.2 $\mu$ distortions and PBH seeds for SMBHs
At the largest scales, the primordial power spectrum is strongly constrained by the CMB observations Aghanim et al. (2020) and the Lyman-$\alpha$ forest data Bird et al. (2011). The CMB observations strongly constrain the curvature power spectrum at scales $10^{-4}\mathrm{Mpc}^{-1}\lesssim k\lesssim 1\mathrm{Mpc}^{-1}$. At redshifts $z\lesssim 10^{6}$, energy injections into the primordial plasma cause persisting spectral distortions in the CMB. These distortions are divided into chemical potential $\mu$-type distortions created at early times and Compton $y$-type distortions created at $z\lesssim 5\times 10^{4}$. For a given curvature power spectrum $\mathcal{P}_{\zeta}(k)$ the spectral distortions are Chluba et al. (2012b); Chluba and Grin (2013)
$X=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{k}\mathcal{P}_{\zeta}(k)W_{X}(k)$ (36)
with $X=\mu,y$ where $k_{\min}=1$ $\mathrm{Mpc}^{-1}$ and the window functions can be approximated by
$W_{\mu}(k)=2.2\left[e^{-\frac{(\hat{k}/1360)^{2}}{1+(\hat{k}/260)^{0.6}+\hat{k}/340}}-e^{-(\hat{k}/32)^{2}}\right],$ (37)
$W_{y}(k)=0.4e^{-(\hat{k}/32)^{2}}$ (38)
with $\hat{k}=k/\left(1\mathrm{Mpc}^{-1}\right)$. The COBE/Firas observations constrain the $\mu$ distortions as $\mu\leq 4.7\times 10^{-5}$ Bianchini and Fabbian (2022) and $y\leq 1.5\times 10^{-5}$ at $95\%$ confidence level Fixsen et al. (1996). The situation is summarised in Fig. 5. Including NGs modifies only the tails of the constraints and not the most constrained region Sharma et al. (2024).
Consequently, the amplitude of the power spectrum in the range of scales related to the FIRAS experiment is constrained to be at most $A=10^{-5}$.
We focus on the threshold statistics approach and we compute the mass fraction $\beta$ as a function of the parameters $f_{\rm NL}$ and $g_{\rm NL}$, assuming the power series ansatz as in Eq. 5. The power series expansion holds until we are in the perturbative regime, that is, the terms are ordered hierarchically with the higher orders being typically smaller than the lower ones. For a narrow power spectrum, this translates into a constraint for the amplitude of the power spectrum, $(3/5)\left|f_{\mathrm{NL}}\right|A^{1/2}\ll 1$, and $(9/25)\left|g_{\mathrm{NL}}\right|A\ll 1$.
Such large positive NGs can lead to the production of a non-negligible population of massive PBHs that can seed SMBHs.
Figure 6: Left panel: Computation of the mass fraction $\beta$ changing the NG parameters $f_{\rm NL}$ and $g_{\rm NL}$ with two benchmark cases for the log-normal power spectrum with an amplitude fixed to be $A=10^{-5}$. Right panel: Computation of the PBH abundance $f_{\rm PBH}$ changing the scale of the peak $k_{*}$ and the NG parameter $g_{\rm NL}$. We fix $\Delta=1$, $A=10^{-5}$ and $f_{\rm NL}=1$.
Indeed, the origin of SMBHs presents a significant challenge in the field of astrophysics. Although it is well known that they occupy the centres of most galaxies Kormendy and Richstone (1995); Magorrian et al. (1998); Richstone et al. (1998), the processes leading to their formation remain unclear.
In the context of PBHs, even a minimal presence of massive PBHs within the mass range $10^{3}-10^{6}\,M_{\odot}$ could potentially seed for SMBHs Duechting (2004); Kohri et al. (2014); Bernal et al. (2018). As a naive computation to estimate the necessary abundance of these seeds, we follow Refs. Vaskonen and Veermäe (2021); Serpico et al. (2020). Considering that SMBHs make up approximately $\mathcal{O}(10^{-4})$ of the stellar mass in their host galaxies Reines and Volonteri (2015), and that stars account for roughly $\mathcal{O}(10^{-2})$ of the total cosmic matter content Fukugita et al. (1998), we deduce that the overall SMBH density is about $10^{6}$ times less than the dark matter density.
The primordial seeds’ abundance can then be expressed as
$f_{\mathrm{PBH}}\sim\mathcal{O}(10^{-6})\times\left\langle M_{\mathrm{PBH}}\right\rangle/M_{\rm SMBH}\,,$ (39)
with $\left\langle M_{\mathrm{PBH}}\right\rangle>10^{3}M_{\odot}$. Nevertheless, the scales related to the formation of SMBHs are strongly constrained by the analysis of the CMB spectral distortion by the FIRAS collaboration Fixsen et al. (1996); Chluba et al. (2012a, b); Chluba and Grin (2013); Bianchini and Fabbian (2022), and in the gaussian approximation, the corresponding abundance of PBHs is too small to furnish a primordial origin for Supermassive black holes seeds.
As shown in the left panel of Fig. 6, in agreement with the literature Ünal et al. (2021); Nakama et al. (2018); Hooper et al. (2024); Byrnes et al. (2024), quadratic primordial NGs are not enough to produce a sizeable amount of PBHs to be the seed of SMBHs. As a consequence, differently from Ref. Byrnes et al. (2